Dear sir,
Please help me in giving the solution to the foll.problem as I am a chemical engineer in doing simulation softwares
Please reply how to determine the value of THETTA in the foll.problem
THETTA - SIN(THETTA) = 3.4545
I had tried by using Taylor Series but i did't get the correct answer.
Please Solve that Problem and tell me any other simple method is there.
Regards,
Suresh.V

Found 2 solutions by longjonsilver, khwang:
Answer by longjonsilver(2297)

from my rusty memory, i would suggest that you need to use an iterative process (numerical method) to solve this.
Hopefully, you will have covered these on your course. If not, re-submit this for someone better to reply with an analytical solution.
Jon.

Answer by khwang(438)

For this kind of question there is no correct answwer, since we only can
have approximated answers.(dependng how accuracy is required.)
As I showed you to use sin x = x - x^3/6
or x - x^3/6 + x^5/150
to get some values.
If you use Newton's iteation method ,let f(x) = x -sin x - 3.4545
xn+1 = xn - f(xn)/f'(xn) by choosing some xo(say 1) to start.
Whatever methods you use , there is no correct answer.
Kenny